OpenStudy (anonymous):

Solve four and seven-eighths divided by three. one and five-eighths eight-thirteenths two and five twenty-fourths one and one-third

OpenStudy (anonymous):

can you use a calculator?

OpenStudy (anonymous):

Uh me?

OpenStudy (anonymous):

yea are you not allowed to use one? did your teacher specify?

OpenStudy (anonymous):

well 4 7/8 divided by 3 = 13/8 1 5/8 divided by 3 = 13/24 8/13 divided by 3 = 8/39 2 5/24 divided by 3 = 53/72

OpenStudy (anonymous):

so the answer is....

OpenStudy (anonymous):

wait are all those seperate is this question multiple choice..?

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

its 1 and 5/8

OpenStudy (anonymous):

i didnt think it was multiple choice.

OpenStudy (anonymous):

are you sure that's the answer?

OpenStudy (anonymous):

since its multiple choice yea im pretty sure, i got 13/8 which is the improper form of 1 5/8, which is the answer given and since you cant reduce it anymore then yea i would go with that